OpenStudy (anonymous):

find the difference simplify 8s/ s^2-14s+49 - 56/ S^2-14s+49

OpenStudy (anonymous):

\[\frac{ 8s }{ s ^{2}-14s+49 }-\frac{ 56 }{ s ^{2}-14s+49 }\]

OpenStudy (tkhunny):

Why is that hard? The denominators are the same already. Just combine the numerators. No?

OpenStudy (anonymous):

i get confused on the simplify and plus im learning off a computer so its not that easy to learn

OpenStudy (tkhunny):

Fair enough. First, \(\dfrac{a}{b} + \dfrac{c}{b} = \dfrac{a+c}{b}\). It's the way fractions work. How do we apply this to your expression?

OpenStudy (anonymous):

\[\frac{ 8s-56 }{ s ^{2}-14s+49 }\]

OpenStudy (tkhunny):

That's it. Now, start factoring.